Download
2e34 meeting and mcc news n.
Skip this Video
Loading SlideShow in 5 Seconds..
2E34 Meeting and MCC News PowerPoint Presentation
Download Presentation
2E34 Meeting and MCC News

2E34 Meeting and MCC News

86 Vues Download Presentation
Télécharger la présentation

2E34 Meeting and MCC News

- - - - - - - - - - - - - - - - - - - - - - - - - - - E N D - - - - - - - - - - - - - - - - - - - - - - - - - - -
Presentation Transcript

  1. 2E34 Meeting and MCC News IR2 Vacuum Studies (Mike Sullivan) Plans for Lattice and Optics Work (Yunhai Cai)

  2. Typical event pressures Detector Forward Q2 NEG Support tube end bellows 1.9 96 LER HER 1540 630 0 1.9 0 Backward Q2 NEG 0.1 LER frangible link NEG pump M.Sullivan

  3. Helium hosing layout Detector Support tube end bellows LER HER Known small vacuum leak NEG pump M.Sullivan

  4. Beam Currents during He Flow Experiments Vacuum Spike Aborts

  5. Signal from the known small leak about 20-25 m in the upstream LER The flow meter says 2 SCFH for one tube RGA He signal appeared ~5 min after He flow started M.Sullivan

  6. He flow turned on at 17:00. No signal yet. The two ion pumps 8020 and 3044 were turned off M.Sullivan

  7. He flow still on No signal when we get a vacuum spike at 17:37 This spike reads 150 nTorr at 3027 M.Sullivan

  8. He flow still on First He signal at 17:50 M.Sullivan

  9. He signal getting stronger 10 min after first sign M.Sullivan

  10. He signal getting stronger 20 min after first sign Some vacuum activity also seen on the gauges M.Sullivan

  11. He turned off at about 18:20, at left edge of plot The two ion pumps 8020 and 3044 were turned on where the Argon trace drops A spike is seen in the He trace when we get a vacuum spike at 18:50 M.Sullivan

  12. Status and plans • More external leak checking • We need to try to understand the signal we have seen • Presently installing a controlled leak at the 3044 ion pump location. We are using the pump out port located there. • Concentrate on improving the analysis that pinpoints the z location • More events to H. Nicholson of BaBar • Work on the IR vacuum model to see if we can get a z location from it • Study and perhaps improve B. Peterson’s analysis • Hooked up one of the BPMs at 3024 to a high speed cable – did this yesterday – nothing seen yet • Review the borescope videos again – still to do • More… M.Sullivan

  13. How To Improve the Specific Luminosity • Closer to half integer Beam size at the interaction point: • Solenoid compensation • Orthogonal IP knobs • Lower beta function • Orthogonal waist knob At high currents: • Vertical steps in the LER • Solenoid compensation • Vertical dispersion in arcs • Coupling in arcs • e1 may have to be increased • Fast damping time is better Y.Cai

  14. MAC recommendations • Develop a focused and schedule-oriented approach to achieving a high quality optics model and lattice correction capability. • Develop the machinery to allow rapid iteration between measurement and correction. • Aggressively pursue lattice correction in both rings to drive the actual lattices to the design. • Pursue model-based coupling analysis and correction. Y.Cai

  15. Milestones • March 31, 2006 • Measure machine optics and build a linear online model within a shift • Reduce beta beating down to 30% in both machine without losing luminosity • Stabilize the orbit within 200 microns in both machines • Test coupling and dispersion corrections in the machine • Calibrate IP knobs against experiment • June 31, 2006 • Update beam-based on model within two hours for both machines • Reduce beta beating down to 15% in both machines without losing luminosity • Stabilize the orbit within 100 microns in both machines • Build all IP knobs within a shift based on the online model • Reduce ey bellow 1.5 nm-rad in both machines without losing luminosity These are not just straight-forward works. There are associated accelerator physics issues to be solved. Y.Cai